ASSIGNMENT OF THE MM- AND SMM-WAVE ROTATIONAL SPECTRA OF RARE ISOTOPOLOGUES OF CYANAMIDE AND THE r_{m}^{(1)} MOLECULAR GEOMETRY OF NH_{2}CN

\underline{ADAM KRASNICKI}, ZBIGNIEW KISIEL, Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, 02-668 Warszawa, Poland; BRENDA P. WINNEWISSER, MANFRED WIN-NEWISSER, Department of Physics, The Ohio State University, Columbus, OH 43210.

The cyanamide molecule is one of the prototype systems for the study of the large–amplitude inversion motion at the nitrogen atom, and it is also a potential astrophysical species. The mm-wave, smm-wave, and the far infrared spectra of the parent and the two principal deuterated isotopic species are now known in considerable detail.^{a}^{,}^{b}

Presently we report an extended analysis of rotational transitions in the 0^{+} and 0^{−} inversion states for 7 rare isotopic species of cyanamide, measured in the 118-650 GHz frequency region on a deuterated sample with natural abundance of carbon and nitrogen. The spectra of five isotopologues: H_{2}N^{13}CN, HDN^{13}CN, D_{2}N^{13}CN, HD^{15}NCN, and HDNC^{15}N, have been assigned for the first time. For D_{2}^{15}NCN and D_{2}NC^{15}N the knowledge of the rotational spectrum has also been considerably improved relative to preceding work.^{c}^{,}^{d}

The availability of spectroscopic constants for 12 different isotopic species allowed determination of the complete r_{m}^{(1)} and r_{m}^{(1L)} geometries^{e} of cyanamide, providing direct experimental information on the pyramidal nature of the NH_{2} group in cyanamide and on the nonlinearity of the NCN segment. The new experimental geometry is compared with results of ab initio calculations.

aG. Moruzzi et al., J. Mol. Spectrosc., 190, 353-364 (1998).

bZ. Kisiel et al., 63^{rd} OSU International Symposium on Molecular Spectroscopy, WK08, 2008.

cJ. K. Tyler et al., J. Mol. Spectrosc., 43, 248-261 (1972).

dR. D. Brown et al., J. Mol. Spectrosc., 114, 257-273 (1985).

eJ. K. G. Watson et al., J. Mol. Spectrosc., 196, 102-119 (1999).